OpenStudy (anonymous):

Select the equations that are parallel and perpendicular to y = one over fourx + 7 and that pass through the point (−8, 9).

OpenStudy (anonymous):

y = 1/4x + 7 can you tell me the slope of this equation ?

OpenStudy (anonymous):

In y = mx + b form, the slope is in the m position. So what is the slope of y = 1/4x + 7 ?

OpenStudy (anonymous):

let me know when you are ready to participate

OpenStudy (anonymous):

One quarter

OpenStudy (anonymous):

correct. The slope is 1/4. A parallel line will have the same slope. Now we will use y = mx + b (-8,9) x = -8 and y = 9 slope(m) = 1/4 now we sub and solve for b 9 = 1/4(-8) + b 9 = -2 + b 9 + 2 = b 11 = b therefore, our parallel line is : y = 1/4x + 11 perpendicular lines have negative reciprocal slopes. All that means is " flip " the slope and change the sign. So the negative reciprocal of the slope given, 1/4, is -4/1, or just -4. now we use y = mx + b and solve for b (-8,9) x = -8 and y = 9 slope(m) = -4 9 = -4(-8) + b 9 = 32 + b 9 - 32 = b - 23 = b so your perpendicular equation is : y = -4x - 23 any questions ?

OpenStudy (anonymous):

Nope, thank you :)

OpenStudy (anonymous):

anytime :)